OpenStudy (anonymous):

Sin Alpha = -7/25 and Cot Beta = -8/15. Neither Alpha nor Beta is in Quadrant 4. Find cos (alpha +Beta)

Directrix (directrix):

I think the cosine of a sum of two angles will be needed.

OpenStudy (anonymous):

Right, you can find Sin Beta with (-7/25)^2 + Cos^2(alpha) =1

OpenStudy (anonymous):

But my question is how do you convert Cot Beta into Sin and Cos

Directrix (directrix):

Yes, you can.

OpenStudy (anonymous):

-8/15 = Cos/Sin..?

OpenStudy (freckles):

you can make a right triangle \[\cot(\beta)=\frac{a}{o} \]